OpenStudy (anonymous):

how do simplify (2+4i) + (4-i) ?

OpenStudy (anonymous):

combine like terms... 2+4 +4i -i

OpenStudy (anonymous):

6+3i

OpenStudy (anonymous):

add the real part 2+ 4 = 6 add the imaginary part 4i +(-i) = 3i giving 6 + 3i

OpenStudy (amistre64):

+1 * n = n -1 * n = -n so a +(...) is just a useless waste of the paranthesis